Universität Stuttgart (STU)

Institut für Informatik Stuttgart Germany

Researchers Irene Stahl, Birgit Tausend.

ILP Areas Predicate Invention, Declarative bias.

Description of Research The research of the Stuttgart’s Machine Learning Group started with work on inverse resolution (IR) [13, 14] and analogical reasoning [1]. Both areas of research have been pursued further, the first with emphasis on the properties of IR-operators [2] and techniques to invent new predicates [15], the second leading to the idea of an abstract representation of the hypothesis space [9]. In [4, 5] concepts of both have been integrated in an efficient ILP system.

The two main foci of recent research in ILP at Stuttgart are predicate invention (PI) and frameworks for representing bias. PI is a method for recovering from an overly strong bias by extending the vocabulary of the hypothesis language. If the learning task fails in the given biased target language, new predicates might be introduced to shift the bias. A survey of approaches to PI is given in [6]. The work in Stuttgart concentrates on the exploration of decision criteria for PI and its usefulness as bias shift operation [7, 8]. Currently, methods for determining the structure and definition of the necessary new predicates are investigated and developed.

All these topics depend strongly on the language bias the new predicates shall weaken. The unifying framework for declaratively representing biases developed in Stuttgart [9, 11, 12] helps to investigate PI-methods for different language restrictions. Furthermore, this framework allows to describe generalisation and specialisation operators with respect to the bias, and to formulate bias shift operations [10]. The unifying representation facilitates the comparison of biases commonly used in ILP, and helps to classify systems according to the strength of their bias.

Recent Publications

[1] S. Bell, B. Tausend. Analogical reasoning for logic programming. Proceedings of the International Workshop on ILP, 1991.

[2] B. Jung. On inverting generality relations. Proceedings of the 3rd International Workshop on ILP, J. Stefan Institute, Technical Report IJS-DP-6707,1993.

[3] U. Knoll, G. Nakhaeizadeh, B. Tausend. Cost-sensitive pruning of decision trees. Proc. European Conf. on Machine Learning, ECML-94, Catania. Springer-Verlag, 1994.

[4] I. Stahl, B. Tausend, R. Wirth. General-to-specific learning of Horn clauses from positive examples. Proceedings of the CompEuro92, the Hague, the Netherlands, 1992.

[5] I. Stahl, B. Tausend, R. Wirth. Two methods for improving inductive logic programming systems. Proc. European Conf. on Machine Learning, ECML-93, Vienna. Springer-Verlag, 1993.

[6] I. Stahl. Predicate invention in ILP – An Overview. Proc. European Conf. on Machine Learning, ECML-93, Vienna. Springer-Verlag, 1993.

[7] I. Stahl. Properties of inductive logic programming in function-free Horn logic. Proc. European Conf. on Machine Learning, ECML-94, Catania. Springer-Verlag, 1994.

[8] I. Stahl. On the utility of predicate invention in inductive logic programming. Proc. European Conf. on Machine Learning, ECML-94, Catania. Springer-Verlag, 1994.

[9] B. Tausend. Using and adapting schemes for the induction of Horn clauses. Proc. of the ECAI Workshop Logical Approaches to Machine Learning, Vienna, 1992.

[10] B. Tausend. A unifying representation for language restrictions. Proceedings of the 3rd International Workshop on ILP, J. Stefan Institute, Technical Report IJS-DP-6707, 1993.

[11] B. Tausend. Biases and their effects in inductive logic programming. Proc. European Conf. on Machine Learning, ECML-94, Catania. Springer-Verlag, 1994.

[12] B. Tausend. Representing biases for inductive logic programming. Proc. European Conf. on Machine Learning, ECML-94, Catania. Springer-Verlag, 1994.

[13] R. Wirth. Learning by failure to prove. Proceedings of the Third European Working Session on Learning, 1988.

[14] R. Wirth. Completing logic programs by inverse resolution. Proceedings of the Fourth European Working Session on Learning, 1989.

[15] R. Wirth, P. O’Rorke. Constraints on predicate invention. In Proceedings of the 8th International Workshop on ML, Morgan Kaufmann, 1991.

Contact Persons Irene Stahl, Birgit Tausend Institut für Informatik Universität Stuttgart Breitwiesenstr. 20-22 D-70565 Stuttgart Germany Phone: +49-711-7816435 Fax: +49-711-7801045 E-mail: {stahl, tausend}@informatik.uni-stuttgart.de